OpenStudy (anonymous):

solve for x 3(2x-6)=(x-14)/2

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

3(2x-6)=(x-14)/2 6x-18=(x-14)/2 2(6x-18)=x-14 12x-36=x-14 11x=22 x=2